one number is 6 less than another. Three times the smaller number is 2 more than the larger. find the numbers.
Found 2 solutions by bipinjha, Maths68:
Answer by bipinjha(3)

let the greater number be x
so, smaller num is x-6
again from question
3(x-6)=2
or, x-6=2/3
or, x=2/3+6
so smaller num is 2/3
and greater num is 20/3

Answer by Maths68(1474)

. find the numbers.
Let
Smaller Number = x
Larger Number = y
one number is 6 less than another
x=y-6........(1)
Three times the smaller number is 2 more than the larger.
3x=y+2........(2)
Put the value of x from (1) to (2)
3(y-6)=y+2........(2)
3y-18=y+2
3y-y=2+18
2y=20
y=20/2
y=10
Put the value of y in (1)
x=y-6
x=10-6
x=4
Smaller Number = x = 4
Larger Number = y = 10